# Interactive and Hybrid Imitation Learning: Provably Beating Behavior Cloning

### v2

@@ Appendix C Proof for Warm-Stagger
Lemma 13.
Let $\mathcal{B}_{\text{bc}}:=\{\pi\in\mathcal{B}:\pi(s)=\pi^{E}(s)\ \forall s\in D_{\text{off}}\}$ be the set of policies that agree with the expert on the offline dataset of $N_{\mathrm{off}}$ expert trajectories. Assume the expert $\pi^{E}$ is deterministic and realizable. Then, with probability at least $1-\delta$, for all $\pi^{i}$ generated in Warm-Stagger (Algorithm 2), it holds that:
$D_{H}^{2}(P_{\pi^{i}},P_{\pi^{E}})\leq O\left(\frac{\log(\tilde{B}/\delta)}{N_{\mathrm{off}}}\right).$
Furthermore, the returned policy $\hat{\pi}$ satisfies:
$D_{H}^{2}(P_{\hat{\pi}},P_{\pi^{E}})\leq O\left(\frac{H\log(\tilde{B}/\delta)}{N_{\mathrm{off}}}\right).$
Proof.
Let $\mathcal{B}_{\text{bc}}$ denote the realizable class of policies that agree with $\pi^{E}$ on the offline dataset. By the MLE generalization bound [foster2024behavior], for any single policy $\pi\in\mathcal{B}_{\text{bc}}$, we have with probability at least $1-\delta$:
$D_{H}^{2}(P_{\pi},P_{\pi^{E}})\leq O\left(\frac{\log(B/\delta)}{N_{\mathrm{off}}}\right).$
However, in Warm-Stagger, policies $\pi^{i}$ are not selected directly from $\mathcal{B}_{\text{bc}}$, but are instead mixtures over $\mathcal{B}_{\text{bc}}$ at each time step. A full trajectory is therefore sampled by first choosing a policy $\pi^{h}$ at each step $h\in[H]$, inducing an effective policy over sequences from $\tilde{\mathcal{B}}_{\text{bc}}$. Therefore, by appling $\tilde{\mathcal{B}}$ for Theorem 2, the BC guarantee holds for all $\pi\in\tilde{\mathcal{B}}$, with a factor of $\log|\tilde{\mathcal{B}}|\leq H\log B$ in the worst case instead of $\log B$. Therefore, by Lemma 12, each $\pi^{i}$ can be viewed as first-step mixing and satisfies:
$D_{H}^{2}(P_{\pi^{i}},P_{\pi^{E}})\leq O\left(\frac{H\log(\tilde{B}/\delta)}{N_{\mathrm{off}}}\right).$
This remains a worst-case bound. If the per-step base policies are drawn from a factored structure (e.g., $\pi=(\pi_{1},\dots,\pi_{H})$ with $\pi_{h}\in\mathcal{B}_{\text{bc}}$), and the support is shared across steps, the effective covering number can be much smaller, reducing the overhead back to $\log B$.
Finally, the returned policy $\hat{\pi}$ is a first-step mixing of $\{\pi_{1},\dots,\pi_{N}\}$, and thus the same Hellinger bound carries over by convexity. ∎
Theorem 14 (Theorem 5 Restated).
If Algorithm 2 is run with a deterministic expert policy $\pi^{E}$, an MDP $\mathcal{M}$ such that $(\mathcal{M},\pi^{E})$ is $\mu$-recoverable, a policy class $\mathcal{B}$ such that realizability holds, and the online learning oracle $\mathcal{O}$ set as the exponential weight algorithm, then it returns $\hat{\pi}$ such that, with probability at least $1-\delta$,
$J(\hat{\pi})-J(\pi^{\mathrm{E}})\leq O\left(\min\left(\frac{R\log(\tilde{B}/\delta)}{N_{\mathrm{off}}},\frac{\mu H\log(B_{\text{bc}}/\delta)}{N_{on}}\right)\right),$
Proof.
By Lemma 13, with probability at least $1-\delta/2$, for all $\pi\in\{\pi_{1},\cdots,\pi_{N}\}$, it satisfies that $D_{H}^{2}(P_{\pi},P_{\pi^{\mathrm{E}}})\leq O\left(\frac{\log(\tilde{B}/\delta)}{N_{\mathrm{off}}}\right)$. This implies the first-step uniform mixture of $\{\pi_{1},\cdots,\pi_{N}\}$ satisfies
$D_{H}^{2}(P_{\hat{\pi}},P_{\pi^{\mathrm{E}}})=\frac{1}{N}\sum_{i=1}^{N}D_{H}^{2}(P_{\pi^{i}},P_{\pi^{\mathrm{E}}})\leq O\left(\frac{\log(\tilde{B}/\delta)}{N_{\mathrm{off}}}\right)$
By invoking Theorem [foster2024behavior][2.1], we have,
$J(\hat{\pi})-J(\pi^{\mathrm{E}})\leq O\left(\frac{R\log(\tilde{B}/\delta)}{N_{\mathrm{off}}}\right)$
For the second half of the proof, we notice that by definition $\pi^{\mathrm{E}}\in\mathcal{B}_{\text{bc}}$. Then, by applying Theorem 3, with probability at least $1-\delta/2$, the returned $\hat{\pi}$ satisfies
$J(\hat{\pi})-J(\pi^{\mathrm{E}})\leq O\left(\frac{\mu H\log(B_{\text{bc}}/\delta)}{N_{\mathrm{int}}}\right).$
Together, we conclude our proof by applying union bound. ∎
@@ Appendix D Proof for Theorem 7
We formally define the MDP $\mathcal{M}$ from Section 4.2, where the expert policy $\pi^{\mathrm{E}}$ is deterministic and the transition dynamics are time-homogeneous across all steps $h\in[H]$.
@@
  • Action Space $\mathcal{A}$: $A=|\mathcal{A}|$ discrete actions. For each state $s\in{\mathcal{S}}$, there is a unique action $\pi^{\mathrm{E}}(s)$ taken by the expert.
@@
  • Specification of Parameters: In the following proofs, we consider $H\geq 50$, $H\geq\frac{5}{4}\log(10N_{0})$, $A\geq 10H$, $\beta=\frac{8}{H-8}$, $N_{1}\geq 500$, and $N_{1}\geq 160N_{0}$.
We also make the following assumption on the output policy of our learning algorithms (BC, Stagger, and Warm-Stagger). At any stage of learning, denote the set of states that are annotated by the expert by ${\mathcal{S}}_{\text{annotated}}$. Given the annotated (state, expert action) pairs, the learner calls some offline or online learning oracle $\mathbb{A}$ to return a policy $\pi$. We require $\pi$’s behavior as follows:
$\pi(\cdot\mid s)=\begin{cases}\pi^{\mathrm{E}}(\cdot\mid s),&s\in{\mathcal{S}}_{\text{annotated}},\\ (\frac{1}{A},\ldots,\frac{1}{A}),&s\notin{\mathcal{S}}_{\text{annotated}}.\end{cases}$
@@
We next make a simple observation that by the construction of $\mathcal{M}$, the expert policy’s state-visitation distribution in $\mathcal{M}$ is stationary over all time steps:
@@
For MDP $\mathcal{M}$, the expert policy $\pi^{E}$’s visitation distribution at time step $h$, $d_{h}^{\pi^{E}}$ equals $\rho$, for all $h\in[H]$.
@@
The initial distribution gives:
@@
Under $\pi^{\mathrm{E}}$, the induced transition kernel on $\mathbf{E}\cup\mathbf{E^{\prime}}$ satisfies:
@@
Hence $\rho$ is a stationary distribution for the Markov chain induced by $\pi^{E}$ (and $\mathbf{B},\mathbf{B^{\prime}}$ are never reached under $\pi^{E}$). Since $d^{\pi^{E}}_{1}=\rho$ and the dynamics are time-homogeneous, induction gives $\forall h\in[H]$, $d^{\pi^{E}}_{h}=\rho$, i.e.,
$d^{\pi^{E}}_{h}(s)=\begin{cases}\frac{1}{N_{0}(1+\beta)},\;&s\in\mathbf{E},\\ \frac{\beta}{N_{1}(1+\beta)},\;&s\in\mathbf{E^{\prime}},\\ 0,&\text{otherwise}.\end{cases}$
Theorem 15 (Restatement of Theorem 7).
@@ D.1 Lower Bound for Behavior Cloning
Throughout this subsection, we denote by $\mathbf{E}_{\text{annotated}}^{\prime}$ the set of states in $\mathbf{E^{\prime}}$ that are visited and annotated by the expert’s $N_{\mathrm{off}}$ offline trajectories.
@@
expert trajectories, then with probability at least $\frac{1}{2}$, the suboptimality of its returned policy $\hat{\pi}$ is lower bounded by:
@@
In the following, we show that an insufficient number of expert trajectories leads to small $\mathbf{E}_{\text{annotated}}^{\prime}$ and poor coverage on states in $\mathbf{E^{\prime}}$. This, in turn, causes the BC-learned policy to frequently fail to recover and get trapped in the absorbing bad state $\mathbf{b}^{\prime}$, incurring a large suboptimality compared to the expert.
By Lemma 17, when $N_{\mathrm{off}}$ is below the stated threshold, at most $1/10$ of the states in $\mathbf{E^{\prime}}$ is annotated by expert trajectories with probability $\frac{1}{2}$. Suppose we roll out the policy $\tilde{\pi}$ in $\mathcal{M}$; let $\tau$ be the first step such that $s_{\tau}\in\mathbf{E^{\prime}}$.
@@
$\Pr_{\mathcal{M},\tilde{\pi}}(\tau\leq H/5)\geq 0.79,\;\text{and}\;\Pr_{\mathcal{M},\tilde{\pi}}\big(s_{\tau}\notin\mathbf{E}_{\text{annotated}}^{\prime}\mid\tau\leq H/5\big)\geq 0.9.$
We henceforth condition on the event $\{\tau\leq H/5,\;s_{\tau}\notin\mathbf{E}_{\text{annotated}}^{\prime}\}$. Applying Lemma 19, we have, with probability at least $0.9$, $\tilde{\pi}$ takes a wrong action at $s_{\tau}$ and transitions to $\mathbf{b^{\prime}}$, and subsequently never takes the expert recovery action at $\mathbf{b^{\prime}}$. Therefore the trajectory remains in $\mathbf{b^{\prime}}$ from time $H/5$ onward, yielding zero reward for at least $\frac{4H}{5}$ steps.
@@
$|\mathbf{E}_{\text{annotated}}^{\prime}|\;\leq\;\frac{N_{1}}{10}.$
@@
Recall that by Observation 1, $d_{h}^{\pi^{E}}(s)=\frac{\beta}{N_{1}(1+\beta)}=\frac{8}{HN_{1}}$ for $s\in\mathbf{E^{\prime}}$, where $\beta=\frac{8}{H-8}$. Denote $X:=|\mathbf{E^{\prime}}_{\text{annotated}}|$ to be the number of annotated states in $\mathbf{E^{\prime}}$. Consider $N_{\mathrm{off}}$ expert trajectories, each of length $H$, and let $s_{i,h}$ be the state at step $h$ of trajectory $i$. Fix any $s\in\mathbf{E^{\prime}}$. By a union bound over all time indices,
$\Pr(s\in\mathbf{E^{\prime}}_{\text{annotated}})=\Pr\Big(\bigcup_{i=1}^{N_{\mathrm{off}}}\bigcup_{h=1}^{H}\{s_{i,h}=s\}\Big)\leq\sum_{i=1}^{N_{\mathrm{off}}}\sum_{h=1}^{H}\Pr(s_{i,h}=s)=\sum_{i=1}^{N_{\mathrm{off}}}\sum_{h=1}^{H}d_{h}^{\pi^{E}}(s)=\frac{8N_{\mathrm{off}}}{N_{1}}.$
@@
$\Pr(s\in\mathbf{E^{\prime}}_{\text{annotated}})\leq\frac{1}{20}.$
Let $Z_{s}$ be the indicator that $s$ is visited by the expert trajectories, by linearity of expectation,
$\mathbb{E}[X]=\sum_{s\in\mathbf{E^{\prime}}}\mathbb{E}[Z_{s}]=\sum_{s\in\mathbf{E^{\prime}}}\Pr(s\in\mathbf{E^{\prime}}_{\text{annotated}})\leq\frac{N_{1}}{20}.$
Applying Markov’s inequality at the threshold $N_{1}/10$,
@@
For the MDP $\mathcal{M}$, for any policy $\pi$ that agrees with $\pi^{\mathrm{E}}$ on $\mathbf{E}$:
@@
The probability of no $s\in\mathbf{E^{\prime}}$ visit in $\frac{H}{5}$ steps is:
@@
$\Pr_{\mathcal{M},\pi^{\mathrm{E}}}\big(\exists h\in[H/5],s_{t}\in\mathbf{E}^{\prime}\;\big)>1-0.2095=0.79$
@@
Let ${\mathcal{S}}_{\text{unannotated}}\subseteq{\mathcal{S}}$ be the set of unannotated states for a policy $\pi$. Consider MDP $\mathcal{M}$ with action space size $A\geq 10H$.
$\Pr\big(\forall h\in[H]:s_{h}\notin{\mathcal{S}}_{\text{annotated}}\implies\ \pi(s_{h})\neq\pi^{E}(s_{h})\big)\;\geq\;0.9.$
@@
Since for any unannotated state $s$, the learner’s policy selects the expert action with probability exactly $\frac{1}{A}$. Condition on any realization with $|H_{\text{unannotated}}|\leq H$. By union bound,
$\Pr\big(\exists h\in[H]:\ s_{h}\in{\mathcal{S}}_{\text{unannotated}}\text{ and }\pi(s_{h})=\pi^{E}(s_{h})\big)\leq\frac{H}{A}\leq\frac{H}{10H}=0.1.$
Thus
$\Pr\big(\forall h\in[H]:s_{h}\notin{\mathcal{S}}_{\text{annotated}}\implies\ \pi(s_{h})\neq\pi^{E}(s_{h})\big)\;\geq\;0.9.$

### v3

@@ C.2 Proof of Theorem 6
For the first term, we recall from Lemma 17 that with probability $1-\delta/2$, for all policies $\pi^{n}$’s, $J(\pi^{\mathrm{E}})-J(\pi^{n})\leq O\left(\frac{R\log(\tilde{B}/\delta)}{N_{\mathrm{off}}}\right)$. Since $J(\hat{\pi})=\frac{1}{N}\sum_{n=1}^{N}J(\pi^{n})$, we have,
$J(\pi^{\mathrm{E}})-J(\hat{\pi})\leq O\left(\frac{R\log(\tilde{B}/\delta)}{N_{\mathrm{off}}}\right).$
For the second term, we note that by definition $\pi^{\mathrm{E}}\in\mathcal{B}_{\mathrm{bc}}$. Thus, by applying Theorem 3, we conclude that with probability at least $1-\delta/2$, the returned $\hat{\pi}$ satisfies
$J(\hat{\pi})-J(\pi^{\mathrm{E}})\leq O\left(\frac{\mu H\log(B_{\mathrm{bc}}/\delta)}{N_{\mathrm{int}}}\right).$
Together, we conclude our proof by applying a union bound. ∎
@@ Appendix D Proof for Theorem 8
We formally define the MDP $\mathcal{M}$ and the expert policy from Section 4.2, where the expert policy $\pi^{\mathrm{E}}$ is deterministic
@@
  • Action Space $\mathcal{A}$: contains $A=|\mathcal{A}|$ discrete actions. For each state $s\in{\mathcal{S}}$, there is a unique action $\pi^{\mathrm{E}}(s)$ taken by the expert.
@@
  • Specification of Parameters: In the following proofs, we let $H\geq\max(50,\frac{5}{4}\log(10N_{0}))$, $A\geq 10H$, $\beta=\frac{8}{H-8}$, $N_{1}\geq\max(500,160N_{0})$.
We also make the following assumption on the policies produced in our learning algorithms (BC, Stagger, and Warm-Stagger). At any stage of learning, denote the set of states that are annotated by the expert by ${\mathcal{S}}_{\mathrm{annotated}}$. Given the annotated (state, expert action) pairs, the learner calls some offline or online learning oracle $\mathbb{A}$ to obtain a policy $\pi$. We require $\pi$’s behavior as follows:
$\pi(\cdot\mid s)=\begin{cases}\pi^{\mathrm{E}}(\cdot\mid s),&s\in{\mathcal{S}}_{\mathrm{annotated}},\\ (\frac{1}{A},\ldots,\frac{1}{A}),&s\notin{\mathcal{S}}_{\mathrm{annotated}}.\end{cases}$
@@
Denote by $d_{h}^{\pi}(s)=\mathbb{P}^{\pi}(s_{h}=s)$ policy $\pi$’s state visitation distribution in $\mathcal{M}$ at step $h$. We next make a simple observation that by the construction of $\mathcal{M}$, the expert policy’s state-visitation distribution in $\mathcal{M}$ is stationary over all steps:
@@
For MDP $\mathcal{M}$, $d_{h}^{\pi^{E}}$ equals $\rho$, for all $h\in[H]$.
@@
Recall the initial state distribution:
@@
Under $\pi^{\mathrm{E}}$, states $\mathbf{b}$ and $\mathbf{b}^{\prime}$ are never reached. And the induced transition kernel on $\mathbf{E}\cup\mathbf{E^{\prime}}$ satisfies:
@@
Hence $\rho$ is a stationary distribution for the Markov chain induced by rolloing out $\pi^{E}$ in $\mathcal{M}$. Since $d^{\pi^{E}}_{1}=\rho$ and the dynamics are time-homogeneous, induction gives that $\forall h\in[H]$, $d^{\pi^{E}}_{h}=\rho$.
The following is the main theorem of this section; setting $N_{0}=\Theta(N_{1})$ (in which case both $N_{0}$ and $N_{1}$ are $\Theta(S)$) gives Theorem 8 in our main paper.
Theorem 19 (Strengthening of Theorem 8).
@@ D.1 Lower Bound for Behavior Cloning
Throughout this subsection, we denote by $\mathbf{E}_{\mathrm{annotated}}^{\prime}$ the set of states in $\mathbf{E^{\prime}}$ that are visited and annotated by the expert’s $N_{\mathrm{off}}$ offline trajectories.
@@
iid expert trajectories, then with probability at least $\frac{1}{2}$, the suboptimality of its returned policy $\hat{\pi}$ is lower bounded by:
@@
In the following, we show that an insufficient number of expert trajectories leads to small $|\mathbf{E}_{\mathrm{annotated}}^{\prime}|$ i.e., poor coverage on $\mathbf{E^{\prime}}$. This, in turn, causes policy $\tilde{\pi}$ to frequently fail to recover and get trapped in the absorbing bad state $\mathbf{b}^{\prime}$, incurring a large suboptimality compared to the expert.
By Lemma 21, when $N_{\mathrm{off}}$ is below the stated threshold, with probability $\frac{1}{2}$,
$|\mathbf{E}_{\mathrm{annotated}}^{\prime}|\leq\frac{1}{10}|\mathbf{E}^{\prime}|$
We henceforth condition on this happening. In this case, recall that our offline learning oracle is such that $\tilde{\pi}$ takes actions uniformly at random in states in $\mathbf{E}^{\prime}\setminus\mathbf{E}_{\mathrm{annotated}}^{\prime}$. Suppose we roll out the policy $\tilde{\pi}$ in $\mathcal{M}$; let $\tau$ be the first step such that $s_{\tau}\in\mathbf{E^{\prime}}$ (and $\tau:=H+1$ if no such step exists).
@@
$\Pr_{\mathcal{M},\tilde{\pi}}(\tau\leq H/5)\geq 0.79,\;\text{and}\;\Pr_{\mathcal{M},\tilde{\pi}}\big(s_{\tau}\notin\mathbf{E}_{\mathrm{annotated}}^{\prime}\mid\tau\leq H/5\big)\geq 0.9.$
We henceforth condition on the event $\{\tau\leq H/5,\;s_{\tau}\notin\mathbf{E}_{\mathrm{annotated}}^{\prime}\}$ when rolling out $\tilde{\pi}$. Applying Lemma 23, we have, with probability at least $0.9$, $\tilde{\pi}$ takes a wrong action at $s_{\tau}$ and transitions to $\mathbf{b^{\prime}}$, and subsequently never takes the expert recovery action at $\mathbf{b^{\prime}}$. Therefore the trajectory remains in $\mathbf{b^{\prime}}$ from step $H/5$ onward, yielding zero reward for at least $\frac{4H}{5}$ steps.
@@
$|\mathbf{E}_{\mathrm{annotated}}^{\prime}|\;\leq\;\frac{N_{1}}{10}.$
@@
Recall that by Observation 1, $d_{h}^{\pi^{E}}(s)=\frac{\beta}{N_{1}(1+\beta)}=\frac{8}{HN_{1}}$ for $s\in\mathbf{E^{\prime}}$, where $\beta=\frac{8}{H-8}$. Denote $X:=|\mathbf{E}_{\mathrm{annotated}}^{\prime}|$ to be the number of annotated states in $\mathbf{E^{\prime}}$. Consider $N_{\mathrm{off}}$ expert trajectories, each of length $H$, and let $s_{i,h}$ be the state at step $h$ of trajectory $i$. Fix any $s\in\mathbf{E^{\prime}}$. By a union bound over all steps and available offline expert trajectories,
$\Pr(s\in\mathbf{E}_{\mathrm{annotated}}^{\prime})=\Pr\Big(\bigcup_{i=1}^{N_{\mathrm{off}}}\bigcup_{h=1}^{H}\{s_{i,h}=s\}\Big)\leq\sum_{i=1}^{N_{\mathrm{off}}}\sum_{h=1}^{H}\Pr(s_{i,h}=s)=\sum_{i=1}^{N_{\mathrm{off}}}\sum_{h=1}^{H}d_{h}^{\pi^{E}}(s)=\frac{8N_{\mathrm{off}}}{N_{1}}.$
@@
$\Pr(s\in\mathbf{E^{\prime}}_{\mathrm{annotated}})\leq\frac{1}{20}.$
Let $Z_{s}$ be the indicator that state $s$ appears in the expert trajectories, by linearity of expectation,
$\mathbb{E}[X]=\sum_{s\in\mathbf{E^{\prime}}}\mathbb{E}[Z_{s}]=\sum_{s\in\mathbf{E^{\prime}}}\Pr(s\in\mathbf{E^{\prime}}_{\mathrm{annotated}})\leq\frac{N_{1}}{20}.$
Applying Markov’s inequality at threshold $N_{1}/10$,
@@
For the MDP $\mathcal{M}$, and any policy $\pi$ that agrees with $\pi^{\mathrm{E}}$ on $\mathbf{E}$:
@@
The probability that no $s\in\mathbf{E^{\prime}}$ is visited in $\frac{H}{5}$ steps is:
@@
$\Pr_{\mathcal{M},\pi^{\mathrm{E}}}\big(\exists h\in[H/5],s_{t}\in\mathbf{E}^{\prime}\;\big)>1-0.2095=0.79\qed$
@@
Consider MDP $\mathcal{M}$ with action space size $A\geq 10H$.
$\Pr_{\mathcal{M},\tilde{\pi}}\big(\forall h\in[H]:s_{h}\in(\mathbf{E}^{\prime}\setminus\mathbf{E}_{\mathrm{annotated}}^{\prime})\cup\mathbf{B^{\prime}}\implies\ a_{h}\neq\pi^{E}(s_{h})\big)\;\geq\;0.9.$
@@
Since for any state in $(\mathbf{E}^{\prime}\setminus\mathbf{E^{\prime}}_{\mathrm{annotated}})\cup\mathbf{B^{\prime}}$, policy $\tilde{\pi}$ selects the expert action with probability exactly $\frac{1}{A}$, thus,
$\Pr_{\mathcal{M},\tilde{\pi}}\big(\exists h\in[H]:\ s_{h}\in(\mathbf{E}^{\prime}\setminus\mathbf{E}_{\mathrm{annotated}}^{\prime})\cup\mathbf{B^{\prime}}\text{ and }a_{h}=\pi^{E}(s_{h})\big)$
$\leq$ $\sum_{h=1}^{H}\Pr_{\mathcal{M},\tilde{\pi}}\big(s_{h}\in(\mathbf{E}^{\prime}\setminus\mathbf{E}_{\mathrm{annotated}}^{\prime})\cup\mathbf{B^{\prime}}\text{ and }a_{h}=\pi^{E}(s_{h})\big)$
$\leq$ $\sum_{h=1}^{H}\Pr_{\mathcal{M},\tilde{\pi}}\big(a_{h}=\pi^{E}(s_{h})\mid s_{h}\in(\mathbf{E}^{\prime}\setminus\mathbf{E}_{\mathrm{annotated}}^{\prime})\cup\mathbf{B^{\prime}}\big)\leq\frac{H}{A}\leq 0.1,$
where we recall that $A\geq 10H$ in our construction. The lemma now follows by taking the complement of the probability above. ∎